### v1

# TopoX: A Suite of Python Packages for Machine Learning on Topological Domains

Abstract
We introduce TopoX, a Python software suite that provides reliable and user-friendly building blocks for computing and machine learning on topological domains that extend graphs: hypergraphs, simplicial, cellular, path and combinatorial complexes. TopoX consists of three packages: TopoNetX facilitates constructing and computing on these domains, including working with nodes, edges and higher-order cells; TopoEmbedX provides methods to embed topological domains into vector spaces, akin to popular graph-based embedding algorithms such as node2vec; TopoModelX is built on top of PyTorch and offers a comprehensive toolbox of higher-order message passing functions for neural networks on topological domains. The extensively documented and unit-tested source code of TopoX is available under MIT license at https://github.com/pyt-team.
Keywords: topological deep learning, topological neural networks, graph neural networks, machine learning, Python packages.
1 Introduction
Deep learning has traditionally focused on processing data within Euclidean domains, as exemplified by the development of convolutional neural networks (CNNs) (Krizhevsky et al., 2012) and recurrent neural networks (RNNs) (Sutskever et al., 2014; Bahdanau et al., 2014). These models have been tailored to handle data supported on regular, grid-like structures present in images or sequences, capitalizing on information inherent within Euclidean domains. To process data supported on more diverse domains, geometric deep learning (GDL) has recently emerged as a powerful paradigm, extending the capabilities of deep learning to non-Euclidean data (Bronstein et al., 2021; Zhou et al., 2020; Cao et al., 2020). GDL leverages principles of geometric regularity, such as symmetries, invariances, or equivariances, introducing suitable inductive biases when processing diverse data domains, including sets, grids, manifolds, and graphs. In this context, Topological Deep Learning (TDL) (Hajij et al., ) represents a new, burgeoning frontier in the field of machine learning. TDL explores the development of models that go beyond traditional graph-based abstractions and provides a principled approach to process data that reside in domains characterized by multi-way relations such as simplicial complexes, cell complexes, and hypergraphs. These extensions allow for the representation of diverse data domains encountered in scientific computations. (Feng et al., 2019; Schaub et al., 2020; Hajij et al., 2020; Schaub et al., 2021; Roddenberry et al., 2021; Giusti et al., 2022; Yang and Isufi, 2023; Barbarossa and Sardellitti, 2020). While there have been significant advancements in the theory of topological deep learning, its practical implementation has faced challenges due to the absence of accessible software libraries that simplify computations and support deep learning models using higher-order structures.
In this paper, we present TopoX, an open-source suite of Python packages designed for machine learning and deep learning operations in topological domains. TopoX is organized into three Python packages: TopoNetX, TopoEmbedX, and TopoModelX. These packages enhance and generalize functionalities found in popular mainstream graph computations and learning tools, enabling them on topological domains. What sets TopoX apart is its abstract general design and the exploitation of the resulting modeling flexibility in the implementation of a broad spectrum of topological domains and TDL models (see Table 1). Further, every domain in TopoNetX offers utilities to work with various components such as nodes, edges, and higher-order cells. TopoNetX also supports computations using incidence matrices, (co)adjacency matrices, and up, down, and Hodge Laplacians. From a representation learning point of view, TopoEmbedX provides methods for embedding topological domains, or parts of these domains, into Euclidean domains. TopoModelX offers a wide range of TDL models based on a comprehensive implementation of higher-order message passing, built on the PyTorch framework (Paszke et al., 2019).
The core objectives of TopoX are as follows: facilitate research in topological domains by providing foundational code to understand concepts, and offer a platform to disseminate algorithms; broaden the accessibility of the field by delivering user-friendly topological learning algorithms to the machine learning community; serve as a learning resource for topological domains and TDL, enriched by diverse examples, notebooks, and visualization capabilities; and provide a unified application programming interface (API) that operates on topological domains. Given that topological spaces generalize most data domains encountered in scientific computations, its unified API offers multiple advantages: it enhances interoperability, streamlines productivity, simplifies learning, and fosters collaboration. By providing a common framework, it reduces maintenance, promotes code portability, and supports parallel computing.
2 Implementation Overview
The TopoNetX package is organized into three main modules: classes, algorithms, and transform. The classes module implements numerous common topological domains, such as SimplicialComplex, CellComplex, and more; inheriting from the abstract class Complex. The algorithms module implements spectral methods, distance computation, and connected component analysis in topological domains. The transform module facilitates conversions between different topological domains. TopoNetX uses Numpy and Scipy backends, and offers toy datasets and examples to facilitate learning and improve understanding.
The TopoEmbedX package supports the learning of representations for all topological domains available in TopoNetX. This package contains the module classes, which implements topological representation learning algorithms that generalize the most popular graph-based representation learning algorithms. These algorithms include DeepCell and Cell2Vec. The TopoEmbedX package has an API inspired by scikit-learn (Pedregosa et al., 2011) and utilizes the KarateClub (Rozemberczki et al., 2020) backends.
TopoModelX is a Python package for topological deep learning, providing efficient tools to implement topological neural networks (TNNs). The package consists of two main modules: base and nn. The base module implements higher-order message passing methods, allowing the construction of general-purpose TNNs using the tensor diagram formalism introduced in Hajij et al. and surveyed in Papillon et al. (2023). The nn module implements various TNNs on popular topological domains, such as simplicial complexes, cell complexes, hypergraphs, and combinatorial complexes. Each implementation includes a corresponding Jupyter notebook tutorial for a user-friendly initiation into TDL. TopoModelX leverages PyTorch and PyG (PyTorch Geometric) backends (Fey and Lenssen, 2019). The TopoModelX package is the outcome of a coding challenge that crowd-sourced the implementations of TDL models et al. (2023).
3 Comparison and Interaction with Other Packages
The libraries most closely related to TopoNetX are NetworkX (Hagberg et al., 2008) and HypernetX (Liu et al., 2021). They facilitate computations on graphs and hypergraphs, respectively. TopoNetX utilizes a similar API to these two libraries to faciliaite rapid adoption of topological domains, such as simplicial complexes, cell complexes, and colored hypergraphs. The XGI package (Landry et al., 2023) offers hypergraph functionalities similar to HypernetX with additional support for simplicial complexes and directed hypergraphs.
The closest package to TopoEmbedX is KarateClub (Rozemberczki et al., 2020), which is a Python package consisting of methods for unsupervised learning on graph-structured data. TopoEmbedX extends the functionality of KarateClub to topological domains supported in TopoNetX.
PyG and DGL (Deep Graph Library) (Wang et al., 2019), which are two popular geometric deep learning packages that support deep learning models on graphs, are closely related to TopoModelX. The DHG (Deep HyperGraph) package (Feng et al., 2019) supports deep learning models defined on hypergraphs.
All three of our packages feature a continuous integration pipeline and are well-tested with code coverage of >=95% per package. This is comparable or better than many related graph-based learning libraries, such as PyG, DGL, XGI, NetworkX, and HyperNetX.
| Comparison between TopoNetX and other Python packages | ||
| Packages | Domains | Operations |
| TopoNetX |
Graphs, colored hypergraphs, simplicial complexes, path complexes, cell complexes, combinatorial complexes |
add_node, add_simplex, add_cell, adjacency_matrix, coadjacency_matrix, incidence_matrix, hodge_laplacian_matrix |
| NetworkX |
Graphs |
add_node, add_edge, adjacency_matrix, incidence_matrix, laplacian_matrix |
| HyperNetX |
Hypergraphs |
add_node, add_edge, adjacency_matrix, incidence_matrix |
| XGI |
Simplicial complexes, hypergraphs, dihypergraph |
add_node, add_edge, adjacency_matrix, boundary_matrix, hodge_laplacian |
| Comparison between TopoEmbedX and other Python packages | ||
| Packages | Domains | Embedding algorithms |
| TopoEmbedX |
Graphs, colored hypergraphs, simplicial complexes, path complexes, cell complexes, combinatorial complexes |
Cell2Vec, DeepCell, CellDiff2Vec, HigherOrderLaplacianEigenMap, HOPE, HOGLEE, |
| Karateclub |
Graphs |
Node2Vec, Graph2Vec, Diff2Vec, GL2Vec, IGE, Role2Vec, GraRep |
| Comparison between TopoModelX and other Python packages | ||
| Packages | Domains | Push-forward operators |
| TopoModelX |
Graphs, (colored) hypergraphs, simplicial complexes, path complexes, cell complexes, combinatorial complexes |
(Higher order) message passing, merge operator, split operator |
| DGL |
Graphs |
Message passing |
| DHG |
Graphs, hypergraphs |
Message passing, hypergraph message passing |
| PyG |
Graphs |
Message passing |
4 Usage: Elementary Examples
TopoNetX provides an user-friendly interface which allows to create a complex in two main steps: first, instantiate the complex; second, add cells to that complex, as shown in the three first lines of the code snippet below. Processing data on a complex requires matrices that describe the (co)adjacency of the incidence relations among cells, as computed below.
The following code snippet shows how TopoEmbedX embeds edges of the Stanford bunny dataset using the Cell2Vec algorithm (Hajij et al., 2020):
The following code snippet shows how to instantiate, and run the forward-pass of a simplicial neural network (SNN) with TopoModelX:
TopoNetX provides a high-level declarative interface. In TopoEmbedX, each embedding algorithm is compatible with all complexes available in TopoNetX. In TopoModelX, TNNs are classified according to the topological domain upon which they are defined. In each package, the directories that contain examples and notebooks offer an abundance of code snippets to assist users in starting their journey with TopoX.
Acknowledgments
M. H. acknowledges support from the National Science Foundation, award DMS-2134231. N. M. acknowledges support from the National Science Foundation, Award DMS-2134241. T. B. acknowledges support from the Engineering and Physical Sciences Research Council [grant EP/X011364/1]. T. K. D. acknowledges support from the National Science Foundation, Award CCF 2049010. N. L. acknowledges support from the Roux Institute and the Harold Alfond Foundation. R. W. acknowledges support from the National Science Foundation, Award DMS-2134178. P. R. acknowledges support from the National Science Foundation, Award IIS-2316496. M. T. S. and F.F. acknowledge funding by the Ministry of Culture and Science (MKW) of the German State of North Rhine-Westphalia (NRW Rückkehrprogramm). M.T.S and J.H. acknowledge funding from the European Union (ERC, HIGH-HOPeS, 101039827). Views and opinions expressed are however those of M. T. S. only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency; neither the European Union nor the granting authority can be held responsible for them. The authors acknowledge the ICML Topological Deep Learning challenge that jump started the development of the TMX part of the software suite.